OpenStudy (anonymous):

find a*b (dot product) if |a|=6 and |b|=5, the angle between a and b is 2pi/3

OpenStudy (anonymous):

\[a.b = |a| * |b| * \cos \theta \]

OpenStudy (anonymous):

\[\cos 2pie/3 = \cos 120 = \cos( 180 - 60) = -\cos 60 = \frac{ -1 }{ 2 }\]

OpenStudy (anonymous):

k then is 6+5*-1/2

OpenStudy (anonymous):

6 * 5 * -1/2

OpenStudy (anonymous):

k

OpenStudy (anonymous):

-15

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

k thank

OpenStudy (anonymous):

Welcome